warren has 40 coins all nickles, dimes, and quarters worth 4.05$. He has seven more nickles then dimes how many quarters.


Lori clames she has 20.00$ in quarters ,half dollars, and one dollar bills. She has twice as many quarters as half dollars and Half as many one dollar biils as Half dollars. Give an argument as to why this is wrong.

Answer by Paul(988)

Let dimes be x
Let nickels be 7+x
Let quarters be y
7+x+x+y=40
y=33-2x

5(7+x)+10(x)+25(33-2x)=405
35+5x+10x+825-50x=405
-35x=-455
x=13
13+7=20
20+13=33, 40-33=7

Hence he has 13 dimes, 20 nickels and 7 quartes

Check: 10(13)+5(20)+25(7)
130+100+175=405
405/100 = $4.05